# On the importance of parallel magnetic-field fluctuations for electromagnetic instabilities in STEP

### v1

@@ Abstract
This paper discusses the importance of parallel perturbations of the magnetic-field in gyrokinetic simulations of electromagnetic instabilities and turbulence at mid-radius in the burning plasma phase of the conceptual high-$\beta$, reactor-scale, tight-aspect-ratio tokamak STEP. Previous studies have revealed the presence of unstable hybrid kinetic ballooning modes (hKBMs) and subdominant microtearing modes (MTMs) at binormal scales approaching the ion Larmor radius. It was found that the hKBM requires the inclusion of parallel magnetic-field perturbations to be linearly unstable. Here, the extent to which the inclusion of fluctuations in the parallel magnetic-field can be relaxed is explored through gyrokinetic simulations. In particular, the frequently used MHD approximation (dropping $\delta\!B_{\parallel}$ and setting the $\nabla B$ drift frequency equal to the curvature drift frequency) is discussed and simulations explore whether this approximation is useful for modelling STEP plasmas. It is shown that the MHD approximation can reproduce some of the linear properties of the full STEP gyrokinetic system, but nonlinear simulations using the MHD approximation result in very different transport states. It is demonstrated that it is difficult to improve the agreement even in conditions that appear to be more compatible with the assumptions that underlie the MHD approximation. Furthermore, it is shown that the sensitivity of STEP to $\delta\!B_{\parallel}$ fluctuations is primarily because the plasma sits close to marginality and it is shown that in slightly more strongly driven conditions the hKBM is unstable without $\delta\!B_{\parallel}.$ Crucially, it is demonstrated that the state of large transport typically predicted by local electromagnetic gyrokinetic simulations of STEP plasmas is not solely due to $\delta\!B_{\parallel}$ physics.
@@ 1 Introduction
Reduced core-plasma transport models are essential for the integrated modelling of tokamak plasma scenarios, since first principles approaches are generally too computationally expensive to evaluate over confinement or resistive discharge timescales. Conventional-aspect-ratio devices tend to be dominated by electrostatic microturbulence driven by instabilities, such as modes driven by the ion-temperature gradient (ITG), modes driven by the electron-temperature gradient (ETG), and trapped-electron-driven modes (TEMs), and several reduced-complexity models have been developed to model electrostatic microturbulence (see e.g. [4, 5]). Such reduced models have been quite successful at predicting turbulent transport in conventional-aspect-ratio present-day devices at low-to-modest $\beta$ [where $\beta=2\mu_{0}p/B^{2}$ is the ratio of the total plasma pressure $p$ to the magnetic-field energy density $B^{2}/(2\mu_{0})$, with $B$ the magnetic-field strength and $\mu_{0}$ the permeability of free space], where the turbulence is predominantly electrostatic: see e.g. [6]. However, in STs, the confinement and turbulent transport, reviewed in [7], are both quantitatively and qualitatively different to that in conventional-aspect-ratio devices. ST geometry has favourable stability properties that allow access to higher-$\beta$ regimes where the turbulence is more electromagnetic in character and the dominant unstable modes including kinetic ballooning modes (KBMs) and microtearing modes (MTMs). These electromagnetic modes, less well understood than their electrostatic counterparts, are not well captured in the most advanced reduced core transport models; transport from electromagnetic turbulence has thus not been reliably captured in scenario modelling for STEP.
@@
Fortunately, most microinstabilities share broad characteristics which are well described by local [^1] All simulations in this paper are local simulations; i.e. simulations are performed in a domain of perpendicular size that is infinitesimal in comparison with the length scale over which the equilibrium varies. Plasma equilibrium gradients are taken to be constant across the simulation domain. linearised $\delta\!f$ gyrokinetics (GK) provided that $k_{\perp}\rho_{s}\sim 1$ and $\rho_{\star s}\equiv\rho_{s}/a\ll 1$ where $\rho_{s}$ is the gyroradius of species $s,$ $a$ is a typical equilibrium length scale, and $k_{\perp}$ is the perpendicular wavenumber of the instability. High-fidelity GK simulations must be exploited: (i) to assess electromagnetic microinstabilities and their associated turbulence in the conceptual power-plant designs developed for STEP; and (ii) to improve the physics basis of reduced models used in scenario design.
@@
In this paper, we concern ourselves with electromagnetic instabilities in STEP, in particular the hKBM [8] that GK simulations indicate to dominate turbulent transport [9] in the conceptual designs of STEP. The inclusion of compressional magnetic fluctuations $\delta\!B_{\parallel}$ was found to be essential in [8] for the hKBM to be unstable in STEP (see also discussion in Section 4). However, $\delta\!B_{\parallel}$ physics is often simply missing from many higher-fidelity (e.g. global, full-$f$) GK codes and modelling tools. Instead, in GK theory and simulations of microinstabilities, it is common practice to neglect $\delta\!B_{\parallel}$, and to compensate for this through an approximation by modifying the magnetic-drift velocity. In this work, we will refer to this compensation as the MHD approximation [^2] In Section 3.2, it will be convenient to describe two separate ‘MHD approximations’, based on two different prescriptions of the magnetic-drift velocity. We will refer to these two different approximations as ‘MHD-1’ and ‘MHD-2’. [^3] In this work, $\delta\!B_{\parallel}$ fluctuations are set to zero in all simulations using the MHD approximation. (see e.g. [10, 11, 12]). The primary focus of this paper is to explore sensitivity of hKBM stability to the inclusion of $\delta\!B_{\parallel}$ and the appropriateness of the MHD approximation in STEP plasmas.
@@ 2 The importance of parallel magnetic-field perturbations
Recent GK analyses for mid-radius of the preferred flat-top operating point, STEP-EC-HD [3], revealed the presence of unstable hybrid KBMs [^4] The term ‘hybrid KBM’ was introduced in [8] to describe the dominant instabilities seen in STEP plasmas; typically these modes share attributes associated with KBM, ITG, and TEM. (hKBM) and subdominant MTMs at binormal scales approaching the ion Larmor radius [8, 9] (note that similar MTMs were reported for an earlier proposed variant of STEP [14]). Nonlinear local GK simulations for the same local equilibrium in [9] find that the hKBM turbulence can drive heat fluxes that exceed the available heating power by orders of magnitude in absence of equilibrium flow shear, and that this large transport state is characterised by highly radially extended turbulent eddies. Further simulations in [9] indicate that a transport steady state can be reached if equilibrium flow shear and/or $\beta^{\prime}$ stabilisation are sufficient at the flat-top to regulate the turbulence. At lower $\beta^{\prime}$ and/or in the absence of equilibrium flow shear, however, hKBMs drive very large turbulent transport in all channels that may challenge access through lower $\beta^{\prime}$ states to the burning flat top.
A natural extension to nonlinear simulations in [9] is to attempt higher-fidelity global gyrokinetic simulations for conditions where local GK predicts large turbulent fluxes. $\delta\!B_{\parallel}$, indicated by local GK to be essential for the hKBM to be unstable in STEP-EC-HD [8], is unfortunately neglected in most global GK codes. To progress towards a global description of hKBM-driven turbulence, we must both (i) implement $\delta\!B_{\parallel}$ in a global-capable code; [^5] Exciting progress is being made in this direction (see e.g. [mishchenko_borchardt_hatzky_kleiber_könies_nührenberg_xanthopoulos_roberg-clark_plunk_2023]), and there are several parallel efforts working towards implementing $\delta\!B_{\parallel}$ globally in both Eulerian codes such as GENE [13] and Particle-In-Cell codes such as ORB5 [16]. and (ii) strive to better understand the sensitive dependence of the hybrid mode on $\delta\!B_{\parallel}.$ This paper focuses on the latter.
@@ 2.1 The MHD approximation for the magnetic-drift velocity
As remarked in Section 1, in GK theory and simulations of microinstabilities it is common to neglect the (typically destabilising[17, 18]) [^6] It was demonstrated in [18] that neglecting $\delta\!B_{\parallel}$ typically stabilises MHD-like modes because this omits the role of $\delta\!B_{\parallel}$ in the “magnetic compression” term in $\delta\!W$ (the potential energy of the plasma). If $\delta\!B_{\parallel}$ is ignored then $\delta\!W$ is minimised incorrectly which has a stabilising effect on pressure-driven modes. parallel magnetic perturbation $\delta\!B_{\parallel}$ at low $\beta,$ and to compensate for this by artificially adding a correction term proportional to $\nabla p$ to the magnetic drift velocity (see e.g. [10, 19, 12] and references therein). This idea of modifying the magnetic drifts, hereinafter referred to as the MHD approximation, is recommended in many GK codes on the neglect of $\delta\!B_{\parallel}$. The merits, or lack thereof, of the MHD approximation have been tested in different GK simulations by various authors (see e.g. [17], [12]). Whilst proper treatment of $\delta B_{\parallel}$ physics is often crucial [17], the cancellation of $\delta\!B_{\parallel}$ and the $\nabla p$ term in the $\nabla{B}$ drift has been demonstrated to be a good approximation when calculating the growth rates of KBMs at low $\beta$ and long wavelengths [12]. Here we explore whether this approximation may be appropriate for use in STEP plasmas [^7] Section 3.3 of [9] noted that the modes with longest perpendicular wavelength are typically expected to drive the most transport in STEP plasmas..
@@
    Testing high turbulent fluxes predicted, but potentially not well resolved, by local GK in some local equilibria is of high priority for STEP design. In the absence of moderate equilibrium flow shear, progress requires global gyrokinetic simulations including profile variation that should better resolve the box-scale streamers observed in local simulations [9]. However, most global-capable codes simply do not include $\delta\!B_{\parallel}$ at present. The MHD approximation has been used in a global code to study electromagnetic turbulence in NSTX plasmas [20].
@@ 3 Linear electromagnetic gyrokinetics and the MHD approximation
We are interested in plasmas that are well described by the GK framework (see e.g., [21, 22]): i.e. we are concerned with fluctuations $\sim e^{\mathrm{i}(\mathbf{k}\cdot\mathbf{r}-\omega t)}$, having characteristic frequency $\omega$ and wavenumbers $k_{\parallel}$ and $k_{\perp}$ parallel and perpendicular to the equilibrium magnetic field direction $\mathbf{b}_{0}=\mathbf{B}_{0}/B_{0},$ that satisfy the standard GK ordering ${\omega}/{\Omega_{s}}\sim{\nu_{ss^{\prime}}}/{\Omega_{s}}\sim{k_{\parallel}}/{k_{\perp}}\sim{q_{s}\delta\phi}/{T_{0s}}\sim{\delta B_{\parallel}}/{B_{0}}\sim{\delta\mathbf{B}_{\perp}}/{B_{0}}\sim\rho_{\star,s}\ll 1,$ where $\Omega_{s}=q_{s}B_{0}/m_{s}$ is the cyclotron frequency of species $s$ with charge $q_{s},$ equilibrium density and temperature $n_{0s}$ and $T_{0s},$ respectively, mass $m_{s}$ and thermal speed $v_{\mathrm{th}s}=\sqrt{2T_{0s}/m_{s}}$, $\nu_{ss^{\prime}}$ is the typical collision frequency. The size of electrostatic perturbations is set by the perturbed electrostatic potential $\delta\phi.$ Electromagnetic perturbations enter GK through $\delta\!B_{\parallel}$ and $\delta\!B_{\perp},$ the fluctuations of the magnetic-field parallel and perpendicular to the equilibrium magnetic-field direction. Electromagnetic effects are most conveniently described in local GK by writing the fluctuating magnetic field $\delta\!\mathbf{B}=\nabla\times(\delta\!\mathbf{A}_{\perp}+\delta\!A_{\parallel}\mathbf{b})\simeq\nabla\times\delta\!\mathbf{A}_{\perp}+\nabla\delta\!A_{\parallel}\times\mathbf{b},$ then relating its parallel component to the fluctuating vector potential $\mathbf{A}$ by $\mathbf{b}\cdot\delta\!\mathbf{B}\equiv\delta\!B_{\parallel}\simeq\mathbf{b}\cdot\nabla\times\delta\!\mathbf{A}_{\perp}.$
@@
where: $b_{s}=k_{\perp}v_{\perp}/\Omega_{s}$; $J_{0}$ and $J_{1}$ are Bessel functions of the first kind that arise from gyroaveraging; the drive frequency $\omega_{\star s}^{T}=\omega_{\star s}\left[1+\eta_{s}\left({v^{2}}/{v^{2}_{\mathrm{th}s}}-3/2\right)\right]$ involves $\eta_{s}=\mathrm{d}\ln T_{0s}/\mathrm{d}n_{0s}$ and the diamagnetic frequency $\omega_{\star s}=({k_{\perp}T_{0s}}/{q_{s}B_{0}})\mathrm{d}\ln n_{0s}/\mathrm{d}r$ is defined in terms of the radial coordinate $r$. The magnetic drift frequency is given by
@@
where the $\nabla B$ drift frequency $\propto\omega_{\nabla B}=\mathbf{k}\cdot\mathbf{b}_{0}\times\nabla B_{0}/B_{0}$, and the curvature drift frequency $\propto\omega_{\kappa}=\mathbf{k}\cdot\mathbf{b}_{0}\times(\mathbf{b}_{0}\cdot\nabla\mathbf{b}_{0}).$ [^8] Note that $\omega_{\kappa}$ and $\omega_{\nabla B}$ [appearing in (2)] are not themselves frequencies. These quantities are related to the curvature and $\nabla B$ drifts, $\Omega_{\kappa s}$ and $\Omega_{\nabla B}$ viz $\Omega_{\kappa s}=[(v_{\mathrm{th}s}^{2})/(2\Omega_{s})]\omega_{\kappa}$ and $\Omega_{\nabla B}=[(v_{\mathrm{th}s}^{2})/(\Omega_{s})]\omega_{\kappa}.$ The equilibrium magnetic field can be written as $\mathbf{B}_{0}=\mathbf{\nabla}\psi\times\mathbf{\nabla}\alpha$, where $\psi$ is poloidal flux, and $\alpha=\xi-q(\psi)\theta-\nu(\psi,\theta)$ is the binormal coordinate, with $\xi$ the toroidal angle, $\theta$ the poloidal angle, and $\nu(\psi,\theta)$ a periodic function of $\theta$ that depends on flux-surface shaping. Perpendicular gradients can be expressed in terms of binormal and radial wavenumbers as $\mathbf{\nabla_{\perp}}=k_{\alpha}\mathbf{\nabla_{\alpha}}+k_{\psi}\mathbf{\nabla_{\psi}}$.
@@
$\frac{\delta\!B_{\parallel}}{B_{0}}=-\frac{\mu_{0}}{B_{0}^{2}}\sum_{s}\int\mathrm{d}\mathbf{v}^{3}\,g_{s}m_{s}v_{\perp}^{2}\,\frac{J_{1}(b_{s})}{b_{s}}$ (5)
@@
$\lim_{b_{s}\to 0}\frac{\delta\!B_{\parallel}}{B_{0}}=-\frac{\mu_{0}}{2B_{0}^{2}}\sum_{s}\int\mathrm{d}\mathbf{v}^{3}\,g_{s}m_{s}v_{\perp}^{2}=-\frac{\mu_{0}}{B_{0}^{2}}\delta P_{\perp}$ (6)

### v2

@@ Abstract
This paper discusses the importance of parallel perturbations of the magnetic-field in gyrokinetic simulations of electromagnetic instabilities and turbulence at mid-radius in the burning plasma phase of the conceptual high-$\beta$, reactor-scale, tight-aspect-ratio tokamak STEP. Previous studies have revealed the presence of unstable hybrid kinetic ballooning modes (hKBMs) and subdominant microtearing modes (MTMs) at binormal scales approaching the ion Larmor radius. In this STEP plasma it was found that the hKBM requires the inclusion of parallel magnetic-field perturbations to be linearly unstable. Here, the extent to which the inclusion of fluctuations in the parallel magnetic-field can be relaxed is explored through gyrokinetic simulations. In particular, the frequently used MHD approximation (dropping $\delta\!B_{\parallel}$ and setting the $\nabla B$ drift frequency equal to the curvature drift frequency) is discussed and simulations explore whether this approximation is useful for modelling STEP plasmas. It is shown that the MHD approximation can reproduce some of the linear properties of the full STEP gyrokinetic system, but is too stable at low $k_{y}$ and nonlinear simulations using the MHD approximation result in very different transport states. It is demonstrated that the MHD approximation is challenged by the high $\beta^{\prime}$ values in STEP, and that the approximation improves considerably at lower $\beta^{\prime}$. Furthermore, it is shown that the sensitivity of STEP to $\delta\!B_{\parallel}$ fluctuations is primarily because the plasma sits close to marginality and it is shown that in slightly more strongly driven conditions the hKBM is unstable without $\delta\!B_{\parallel}.$ Crucially, it is demonstrated that the state of large transport typically predicted by local electromagnetic gyrokinetic simulations of STEP plasmas is not solely due to $\delta\!B_{\parallel}$ physics.
@@ 1 Introduction
Reduced core-plasma transport models are essential for the integrated modelling of tokamak plasma scenarios, since first principles approaches are generally too computationally expensive to evaluate over confinement or resistive discharge timescales. Conventional-aspect-ratio devices tend to be dominated by electrostatic microturbulence driven by instabilities, such as modes driven by the ion-temperature gradient (ITG), modes driven by the electron-temperature gradient (ETG), and trapped-electron-driven modes (TEMs), and several reduced-complexity models have been developed to model electrostatic microturbulence (see e.g. [4, 5]). Such reduced models have been quite successful at predicting turbulent transport in conventional-aspect-ratio present-day devices at low-to-modest $\beta$ [where $\beta=2\mu_{0}p/B^{2}$ is the ratio of the total plasma pressure $p$ to the magnetic-field energy density $B^{2}/(2\mu_{0})$, with $B$ the magnetic-field strength and $\mu_{0}$ the permeability of free space], where the turbulence is predominantly electrostatic: see e.g. [6]. However, in STs, the confinement and turbulent transport, reviewed in [7], are both quantitatively and qualitatively different to that in conventional-aspect-ratio devices. ST geometry has favourable stability properties that allow access to higher-$\beta$ regimes where the turbulence is more electromagnetic in character and the dominant unstable modes include kinetic ballooning modes (KBMs) and microtearing modes (MTMs). These electromagnetic modes, less well understood than their electrostatic counterparts, are not well captured in the most advanced reduced core transport models; transport from electromagnetic turbulence has thus not been reliably captured in scenario modelling for STEP.
@@
Fortunately, most microinstabilities share broad characteristics which are well described by local [^1] All simulations in this paper are local simulations; i.e. simulations are performed in a domain of perpendicular size that is infinitesimal in comparison with the length scale over which the equilibrium varies. Plasma equilibrium gradient length scales are taken to be constant across the simulation domain. linearised $\delta\!f$ gyrokinetics (GK) provided that $k_{\perp}\rho_{s}\sim 1$ and $\rho_{\star s}\equiv\rho_{s}/a\ll 1$ where $\rho_{s}$ is the gyroradius of species $s,$ $a$ is a typical equilibrium length scale, and $k_{\perp}$ is the perpendicular wavenumber of the instability. High-fidelity GK simulations must be exploited: (i) to assess electromagnetic microinstabilities and their associated turbulence in the conceptual power-plant designs developed for STEP; and (ii) to improve the physics basis of reduced models used in scenario design.
@@
In this paper, we concern ourselves with electromagnetic instabilities in STEP, in particular the hybrid-kinetic ballooning mode (hKBM) [8] that GK simulations indicate to dominate turbulent transport [9] in the conceptual designs of STEP. The inclusion of compressional magnetic fluctuations $\delta\!B_{\parallel}$ was found to be essential in [8] for the hKBM to be unstable in STEP (see also discussion in Section 4). However, $\delta\!B_{\parallel}$ physics is often simply missing from many higher-fidelity (e.g. global, full-$f$) GK codes and modelling tools. Instead, in GK theory and simulations of microinstabilities, it is common practice to neglect $\delta\!B_{\parallel}$, and to compensate for this by exploiting an approximate cancellation (valid in certain limits) that relies on the form of $\delta\!B_{\parallel}$ in gyrokinetics and an exact relationship between the equilibrium $\nabla B$ and curvature drifts. In this work, we will refer to this compensation as the MHD approximation [^2] In Section 3.2, we will describe two separate implementations of the ‘MHD approximation’ that have been included in some gyrokinetic codes. We will refer to these two different approximations as ‘MHD-1’ and ‘MHD-2’. In all gyrokinetic simulations using the MHD approximation, $\delta\!B_{\parallel}$ fluctuations are set to zero and a magnetic drift term is dropped to compensate for this. (see e.g. [10, 11, 12]); so called because it relies on (i) an MHD equilibrium force balance constraint on the relationship between the $\nabla B$ and curvature drifts, and (ii) the gyrokinetic form of $\delta\!B_{\parallel}$ which excludes high frequency fast compressional Alfvén waves by enforcing perpendicular force balance on fluctuations $\delta\mathbf{j_{\perp}}\times\mathbf{B}=\mathbf{\nabla_{\perp}\cdot\delta p_{\perp}}$ The primary focus of this paper is to explore sensitivity of hKBM stability to the inclusion of $\delta\!B_{\parallel}$ and the appropriateness of the MHD approximation in STEP plasmas.
@@ 2 The importance of parallel magnetic-field perturbations
Recent GK analyses for mid-radius of the preferred flat-top operating point, STEP-EC-HD [3], revealed the presence of unstable hybrid KBMs [^3] The term ‘hybrid KBM’ was introduced in [8] to describe the dominant instabilities seen in STEP plasmas; typically these modes share attributes associated with KBM, ITG, and TEM. (hKBM) and subdominant MTMs at binormal scales approaching the ion Larmor radius [8, 9] (note that similar MTMs were reported for an earlier proposed variant of STEP [14]). Nonlinear local GK simulations for the same local equilibrium in [9] find that the hKBM turbulence can drive heat fluxes that exceed the available heating power by orders of magnitude in the absence of equilibrium flow shear, and that this large transport state is characterised by highly radially extended turbulent eddies. Further simulations in [9] indicate that a transport steady state can be reached if equilibrium flow shear and/or $\beta^{\prime}$ stabilisation are sufficient at the flat-top to regulate the turbulence. (A transport steady state for a STEP flat-top was recently reported [15] using a physics-based model of transport from hKBM turbulence.) At lower $\beta^{\prime}$ and/or in the absence of equilibrium flow shear, however, hKBMs drive very large turbulent transport in all channels that may challenge access to the burning flat top through lower $\beta^{\prime}$ states.
A natural extension to nonlinear simulations in [9] is to attempt higher-fidelity global gyrokinetic simulations for conditions where local GK predicts large turbulent fluxes. $\delta\!B_{\parallel}$, indicated by local GK to be essential for the hKBM to be unstable in STEP-EC-HD [8], is unfortunately neglected in most global GK codes. To progress towards a global description of hKBM-driven turbulence, we must both (i) implement $\delta\!B_{\parallel}$ in a global-capable code; [^4] Exciting progress is being made in this direction (see e.g. [16, 17, 18]), and there are several parallel efforts working towards implementing $\delta\!B_{\parallel}$ globally in both Eulerian codes such as GENE [13] and Particle-In-Cell codes such as ORB5 [19]. and (ii) strive to better understand the sensitive dependence of the hybrid mode on $\delta\!B_{\parallel}.$ This paper focuses on the latter.
@@ 2.1 The MHD approximation for the magnetic-drift velocity
As remarked in Section 1, in GK theory and simulations of microinstabilities it is common to neglect the (typically destabilising[20, 21]) [^5] It was demonstrated in [21] that neglecting $\delta\!B_{\parallel}$ typically is stabilising for MHD-like modes because this omits the “magnetic compression” term in $\delta\!W$ (the potential energy of the plasma). If $\delta\!B_{\parallel}$ is ignored then $\delta\!W$ is minimised incorrectly which has a stabilising effect on pressure-driven modes. parallel magnetic perturbation $\delta\!B_{\parallel}$ at low $\beta,$ and to compensate for this by artificially adding a correction term proportional to $\nabla p$ to the magnetic drift velocity (see e.g. [10, 22, 12] and references therein). This idea of modifying the magnetic drifts, hereinafter referred to as the MHD approximation, is recommended in many GK codes on the neglect of $\delta\!B_{\parallel}$. The merits, or lack thereof, of the MHD approximation have been tested in different GK simulations by various authors (see e.g. [20], [12]). Whilst proper treatment of $\delta B_{\parallel}$ physics is often crucial [20], the cancellation of $\delta\!B_{\parallel}$ and the $\nabla p$ term in the $\nabla{B}$ drift has been demonstrated to be a good approximation when calculating the growth rates of KBMs at low $\beta$ and long wavelengths [12]. Here we explore whether this approximation may be appropriate for use in STEP plasmas [^6] Section 3.3 of [9] noted that the modes with longest perpendicular wavelength are typically expected to drive the most transport in STEP plasmas..
@@
    Testing high turbulent fluxes predicted, but potentially not well resolved, by local GK in some local equilibria is of high priority for STEP design. Progress requires global gyrokinetic simulations including profile variation that should better resolve the box-scale streamers observed in local simulations [9].However, most global-capable codes simply do not include $\delta\!B_{\parallel}$ at present because of the associated algorithmic complexity. The GK equation is closed by self-consistently solving the gyrokinetic form of Maxwell’s equations to evaluate the perturbed electromagnetic fields. Retaining the perpendicular part of Ampère’s law introduces a coupling between the electrostatic potential and the perpendicular magnetic potential $\delta\!B_{\parallel}.$ Including $\delta\!B_{\parallel}$ consistently requires the solution of a potentially poorly-conditioned system of equations involving $\delta\!\phi$ and $\delta\!B_{\parallel}$, as well as derivatives and gyro-averages of these quantities. In a local code, these derivatives and gyro-averages are trivial to handle via spectral methods. In a global code, such an approach is not possible and one has to instead introduce complex “gyrodisk averages” to obtain a system which is numerically well behaved. (see [17, 18]). The MHD approximation has been used in a global code to study electromagnetic turbulence in NSTX plasmas [23].
@@ 3 Linear electromagnetic gyrokinetics and the MHD approximation
We are interested in plasmas that are well described by the GK framework (see e.g., [24, 25]): i.e. we are concerned with fluctuations $\sim e^{\mathrm{i}(\mathbf{k}\cdot\mathbf{r}-\omega t)}$, having characteristic frequency $\omega$ and wavenumbers $k_{\parallel}$ and $k_{\perp}$ parallel and perpendicular to the equilibrium magnetic field direction $\mathbf{b}=\mathbf{B}/B,$ that satisfy the standard GK ordering ${\omega}/{\Omega_{s}}\sim{\nu_{ss^{\prime}}}/{\Omega_{s}}\sim{k_{\parallel}}/{k_{\perp}}\sim{q_{s}\delta\phi}/{T_{0s}}\sim{\delta B_{\parallel}}/{B}\sim{\delta\mathbf{B}_{\perp}}/{B}\sim\rho_{\star,s}\ll 1,$ where $\Omega_{s}=q_{s}B/m_{s}$ is the cyclotron frequency of species $s$ with charge $q_{s},$ equilibrium density and temperature $n_{0s}$ and $T_{0s},$ respectively, mass $m_{s}$ and thermal speed $v_{\mathrm{th}s}=\sqrt{2T_{0s}/m_{s}}$, $\nu_{ss^{\prime}}$ is the typical collision frequency. The perturbed electrostatic potential is represented by $\delta\phi.$ Electromagnetic perturbations enter GK through $\delta\!B_{\parallel}$ and $\delta\!B_{\perp},$ the fluctuations of the magnetic-field parallel and perpendicular to the equilibrium magnetic-field direction. Electromagnetic effects are most conveniently described in local GK by writing the fluctuating magnetic field $\delta\!\mathbf{B}=\nabla\times(\delta\!\mathbf{A}_{\perp}+\delta\!A_{\parallel}\mathbf{b})\simeq\nabla\times\delta\!\mathbf{A}_{\perp}+\nabla\delta\!A_{\parallel}\times\mathbf{b},$ then relating its parallel component to the fluctuating vector potential $\mathbf{A}$ by $\mathbf{b}\cdot\delta\!\mathbf{B}\equiv\delta\!B_{\parallel}\simeq\mathbf{b}\cdot\nabla\times\delta\!\mathbf{A}_{\perp}.$
@@
where: $b_{s}=k_{\perp}v_{\perp}/\Omega_{s}$; $J_{0}$ and $J_{1}$ are Bessel functions of the first kind that arise from gyroaveraging; the drive frequency $\omega_{\star s}^{T}=\omega_{\star s}\left[1+\eta_{s}\left({v^{2}}/{v^{2}_{\mathrm{th}s}}-3/2\right)\right]$ involves $\eta_{s}=\mathrm{d}\ln T_{0s}/\mathrm{d}n_{0s}$ and the diamagnetic frequency $\omega_{\star s}=({k_{\perp}T_{0s}}/{q_{s}B})\mathrm{d}\ln n_{0s}/\mathrm{d}r$ is defined in terms of the radial coordinate $r$. The magnetic drift frequency is given by
@@
where the $\nabla B$ drift frequency $\propto\omega_{\nabla B}=\mathbf{k}\cdot\mathbf{b}\times\nabla B/B$, and the curvature drift frequency $\propto\omega_{\kappa}=\mathbf{k}\cdot\mathbf{b}\times(\mathbf{b}\cdot\nabla\mathbf{b}).$ [^7] Note that $\omega_{\kappa}$ and $\omega_{\nabla B}$ [appearing in (2)] are not themselves frequencies. The equilibrium magnetic field can be written as $\mathbf{B}_{0}=\mathbf{\nabla}\psi\times\mathbf{\nabla}\alpha$, where $\psi$ is poloidal flux, and $\alpha=\xi-q(\psi)\theta-\nu(\psi,\theta)$ is the binormal coordinate, with $\xi$ the toroidal angle, $\theta$ the poloidal angle, and $\nu(\psi,\theta)$ a periodic function of $\theta$ that depends on flux-surface shaping. Perpendicular gradients can be expressed in terms of binormal and radial wavenumbers as $\mathbf{\nabla_{\perp}}=k_{\alpha}\mathbf{\nabla_{\alpha}}+k_{\psi}\mathbf{\nabla_{\psi}}$.
@@
$\frac{\delta\!B_{\parallel}}{B}=-\frac{\mu_{0}}{B^{2}}\sum_{s}\int\mathrm{d}\mathbf{v}^{3}\,g_{s}m_{s}v_{\perp}^{2}\,\frac{J_{1}(b_{s})}{b_{s}}$ (5)
@@
$\lim_{b_{s}\to 0}\frac{\delta\!B_{\parallel}}{B}=-\frac{\mu_{0}}{2B^{2}}\sum_{s}\int\mathrm{d}\mathbf{v}^{3}\,g_{s}m_{s}v_{\perp}^{2}=-\frac{\mu_{0}}{B^{2}}\delta P_{\perp}$ (6)